# Vision-LSTM: xLSTM as Generic Vision Backbone

### v1

@@ 1 Introduction
Similar to vision adaptions of SSMs [23, 44], ViL exhibits linear computational and memory complexity w.r.t. sequence length which makes it appealing for tasks that benefit from high-resolution images such as medical imaging [8, 18, 38, 41], segmentation [22, 9], or physics simulations [5, 27, 6, 2]. In contrast, ViT’s computational complexity scales quadratically due to the self-attention mechanism, rendering them costly to apply to high-resolution tasks.
@@ 2 Method
Vision-LSTM (ViL) is a generic backbone for computer vision tasks, which is residually built from xLSTM blocks, as visualized in Figure 1. Following ViT [13], ViL first splits an image into non-overlapping patches via a shared linear projection, then adds learnable positional embeddings to each patch token. At the core of ViL are alternating mLSTM blocks, which are fully parallelizable and equipped with a matrix memory and a covariance update rule. Odd mLSTM blocks process patch tokens from top left to bottom right while even blocks go from bottom right to top left.
Figure 1: Schematic overview of Vision-LSTM (ViL). Following ViT [13], an input image is split into patches and linearly projected. Then, a learnable vector is added per position to the patches, producing a sequence of patch tokens. This sequence is then processed by alternating mLSTM blocks where even blocks flip the sequence before and after the mLSTM layer. For classification, ViL averages the first and the last patch followed by a linear classification head.
@@ 3 Experiments
We conduct experiments on ImageNet-1K [12], which contains 1.3M training images and 50K validation images where each image belongs to one of 1000 classes. Our comparison focuses on models that use a sequence modeling backbone and are roughly comparable in parameter counts.
@@
We train ViL models for 800 epochs (tiny, tiny+) or 400 epochs (small, small+, base) on 224x224 resolution with a learning rate of 1e-3 using a cosine decay schedule. Detailed hyperparameters can be found in Appendix 5.
To facilitate a fair comparison against Vision Mamba (Vim) [44], we add additional blocks to our models to match parameter counts for the tiny and small variants (denoted as ViL-T+ and ViL-S+). Note that ViL requires significantly less compute than Vim since ViL traverses the sequence in an alternating way, whereas Vim traverses the sequence twice per block. This holds despite Vim using optimized CUDA kernels, which currently do not exist for mLSTM, and would speed-up ViL even more. We compare runtimes in Appendix A.1 where ViL is up to 69% faster than Vim.
As ViTs are well established in the vision community, they underwent multiple optimization cycles over the years [13, 34, 36, 35, 37, 19]. As this work is the first to apply xLSTM to computer vision, we do not expected to outperform years of hyperparameter tuning of ViTs in all cases. Even so, the results in Table 1 show that ViL shows favorable results to heavily optimized ViT protocols (DeiT, DeiT-II, DeiT-III) on smaller scales where only DeiT-III-S that was trained twice as long as ViL-S performs slightly better. On the “base” scale, ViL outperforms the initial ViT [13] models and achieves comparable results to DeiT [34]. Note that due to high costs of training models at this scale, the hyperparameters for ViL-B are far from optimal. For reference, training ViL-B takes roughly 600 A100 GPU-hours or 19 hours on 32 A100 GPUs.
The performance can be further improved by fine-tuning models in a “long-sequence fine-tuning” setting [44], which fine-tunes the models for 30 epochs with an increased sequence length of 729 by using 50% overlap between consecutive patch tokens.
ViL also demonstrates competitive performance with CNN-based models such as ConvNeXt [24], despite not leveraging the inductive bias inherent to convolutions.
Method #Params ACC
DeiT-T [34] 6M 72.2
DeiT-II-T [36] 6M 73.5
DeiT-III-T (reimpl.) 6M 76.2
Vim-T [44] 7M 76.1
ViL-T 6M 77.3
ViL-T+ 7M 78.1
DeiT-S [34] 22M 79.8
DeiT-II-S [36] 22M 80.7
DeiT-III-S [37] 22M 81.4
Vim-S [44] 26M 80.5
ViL-S 23M 80.7
ViL-S+ 26M 80.9
ViT-B384 [13] 86M 77.9
DeiT-B [34] 86M 81.8
DeiT-II-B [36] 86M 82.7
DeiT-III-B [37] 86M 83.7
ViL-B 88M 81.6*
Method #Params ACC
Long-sequence fine-tuning
DeiT-III-T (reimpl.) 6M 78.3
Vim-T [44] 7M 78.3
ViL-T+ 7M 80.0
DeiT-S [34] 22M 81.4
DeiT-III-S [37] 22M 83.2
Vim-S [44] 26M 81.6
ViL-S+ 26M 82.2
DeiT-B [34] 86M 83.2
DeiT-III-B [37] 86M 84.5
ViL-B 88M 82.7*
CNN-based models
ConvNeXt-T [24] 30M 82.2
ConvNeXt-S [24] 50M 83.1
ConvNeXt-B [24] 89M 83.8
SSM-based models
S4ND-Conv-T [26] 30M 82.2
S4ND-ViT-B [26] 89M 80.4
Table 1: ImageNet-1K classification accuracy. Entries in gray are trained trained for 400 epochs more than ViL. “Long-sequence fine-tuning” fine-tunes models with 50% overlapping patches for 30 epochs. We train DeiT models in the long-sequence fine-tuning setting as DeiT-II has no public checkpoints and DeiT-III only published checkpoints with longer training durations. *due to the large costs of training ViL-B, hyperparameters are far from optimal.
Block Design.
@@
Table 2: Traversing the sequence of patch tokens in multiple directions yields better performance than the plain, uni-directional xLSTM block. However, using multiple directions per block also requires more compute and parameters. Sharing parameters between directions matches #Params but overloads them, leading to performance drops. Using multi-directional designs but alternating between them in consecutive blocks, matches #Params and FLOPS while achieving good performances. While quad-directional designs yields the best performances, they comes with drastically increased runtime. As our current quad-directional implementation is not compatible with torch.compile, it suffers from increased runtime even when alternating between directions. Since quad-directional designs are compute intensive, this study was conducted by training a ViL-T on an ImageNet-1K subset of 100 classes using resolution 128.
@@ 3.1 Classification Design
In order to perform classification with a ViT, the sequence of tokens is pooled into a single token, which is then used as input to a classification head. The most common methods to do this pooling are (i) adding a learnable [CLS] token at the start of the sequence or (ii) averaging all patch tokens to produce an [AVG] token. Whether to use the [CLS] or [AVG] token is typically a hyperparameter, where both variants achieve comparable performances. On the contrary, autoregressive models often require specialized classification designs. For example, Vim [44] requires the [CLS] token to be in the middle of the sequence, suffering heavy performance losses if other classification designs, e.g., an [AVG] token or two [CLS] tokens at start and end of the sequence, are employed. Due to its autoregressive nature, we explore different classification designs for ViL in Table 3. [AVG] averages all patch tokens, “Middle Patch” uses the middle patch token, “Middle [CLS]” uses a [CLS] token in the middle of the sequence, “Bilateral [AVG]” uses the average of the first and last patch token. We find ViL to be relatively robust the classification design where all performances are within 0.6%. We choose “Bilateral [AVG]” over “Middle [CLS]”, as ImageNet-1K is known to have a center bias, i.e., objects are commonly located in the middle of the picture. By using “Bilateral [AVG]” we avoid exploiting this bias, which makes our model more general.
Arch ACC
[AVG] 75.4
Middle Patch 75.9
Middle [CLS] 75.3
Bilateral [AVG] 75.9
Bilateral Concat 76.4
@@
Table 3: How to pool the sequence of patch tokens into a single token for classification? ViL is relatively robust to the classification design. Using the middle patch token and averaging the first and last token (“Bilateral [AVG]”) work similarly well. The good performance of middle patch could be caused by the bias in ImageNet-1K that objects are typically in the center of the image. We therefore choose “Bilateral [AVG]” to avoid exploiting this bias. “Bilateral Concat” concatenates the first and last patch token which introduces additional parameters in the classification head. To save compute, we conduct this study by training a ViL-T on ImageNet-1K for 400 epochs in 192x192 resolution followed by fine-tuning for 20 epochs on 224x224 resolution.
To remain comparable to previous architectures that use a single token as input to the classification head, we average the first and the last patch. To achieve best performances, we suggest to concatenate the two tokens (“Bilateral Concat”) instead of averaging them. This is similar to common practices in self-supervised vision transformers, such as DINOv2 [28], which are trained with two objectives that are attached at the [CLS] and [AVG] token respectively and therefore benefit from concatenating the representation of [CLS] and [AVG] token. This directions has also been explored for vision SSM models [40] where multiple [CLS] tokens are spread out over the sequence and then used as input to a classifier. Similar directions could also improve the performance of ViL.
@@ 4 Conclusion
Motivated by the success of xLSTM in language modeling, we introduced ViL, an adaption of the xLSTM architecture to vision tasks. ViL processes a sequence of patch tokens in alternating fashion. Odd blocks process image patches row-wise from top left to bottom right and even blocks go from bottom right to top left. Our new architecture outperforms SSM-based vision architectures and also optimized ViT models on ImageNet-1K classification. Remarkably, ViL is able to outperform ViT training pipelines, that resulted from years of hyperparameter tuning and transformer improvements, in fair comparisons.
In the future, we see potential in applying ViL when high-resolution images are needed for optimal performance, such as semantic segmentation or medical imaging. In these settings, transformers suffer from high computational costs due to the quadratic complexity of self-attention, which ViL does not have due to its linear complexity. Additionally, improving pre-training schemes (e.g. via self-supervised learning), exploring better hyperparameter settings or transferring techniques from transformers (e.g. LayerScale [35]) are promising directions of ViL.
@@ Acknowledgments
Appendix A Extended Results
A.1 Runtime Comparison of ViL vs Vim

### v2

@@ 1 Introduction
Similar to vision adaptions of SSMs [40, 67, 58], ViL can exhibit linear computational and memory complexity w.r.t. sequence length which makes it appealing for tasks that benefit from high-resolution images such as medical imaging [9, 26, 55, 62], segmentation [37, 10], or physics simulations [6, 43, 7, 2]. In contrast, ViT’s computational complexity scales quadratically due to the self-attention mechanism, rendering them costly to apply to high-resolution tasks.
@@ 2 Method
Vision-LSTM (ViL) is a generic backbone for computer vision tasks, which is residually built from xLSTM blocks, as visualized in Figure 1. Following ViT [18], ViL first splits an image into non-overlapping patches via a shared linear projection, then adds learnable positional embeddings to each patch token. At the core of ViL are alternating mLSTM blocks, which are fully parallelizable and equipped with a matrix memory combined with a covariance update rule. Odd mLSTM blocks process patch tokens from top left to bottom right while even blocks go from bottom right to top left.
Figure 1: Schematic overview of Vision-LSTM (ViL). Following ViT [18], an input image is split into patches and linearly projected. Then, a learnable vector is added per position to the patches, producing a sequence of patch tokens. This sequence is then processed by alternating mLSTM blocks where even blocks flip the sequence before and after the mLSTM layer. For classification, ViL uses the concatenation of the first and the last patch as input to a linear classification head.
@@ 3 Experiments
We pre-train models on ImageNet-1K [17], which contains 1.3M training images and 50K validation images where each image belongs to one of 1000 classes. ViL models are trained for 800 epochs (tiny) or 400 epochs (small, base) on 192x192 resolution with a learning rate of 1e-3 using a cosine decay schedule. Afterwards, the model is fine-tuned on 224x224 resolution for 20 epochs using a learning rate of 1e-5. Detailed hyperparameters can be found in Appendix Table 12.
As ViL is an isotropic architecture (like ViT [18]), we compare it against other isotropic architectures and leave exploration of hierarchical ViL architectures to future work.
Figure 2: Performance overview of ImageNet-1K pre-trained models in relation to pre-training compute. ViL shows strong performances across classification and semantic segmentation tasks.
As ViTs are well established in the vision community, they underwent multiple optimization cycles over the years [18, 51, 53, 52, 54, 27]. Since this work is the first to apply xLSTM to computer vision, we do not expected to outperform years of hyperparameter tuning of ViTs in all cases. Even so, Figure 2 shows an overview of performance metrics in relation to total pre-training compute where ViL performs favorably against heavily optimized transformer protocols (DeiT, DeiT-III) and Vision Mamba (Vim) [67] on ImageNet-1K [17] classification, ADE20K [66] semantic segmentation and VTAB-1K [64] classification. Detailed results are presented in the following sections.
3.1 ImageNet-1K Pre-training
Table 1 relates parameter counts and FLOPS to classification accuracy after pre-training on ImageNet-1K. ViL outperforms heavily optimized ViT protocols and other backbones on the tiny and small scale. While ViL does not outperform all other models on the base scale, evaluations on downstream tasks (as shown later in Table 3 and Table 2) show that ViL-B still learns strong features. Additionally, training a ViL-B for 400 epochs takes roughly 600 A100 GPU-hours or 19 hours on 32 A100 GPUs. Therefore, the hyperparameter configuration is most likely not optimal as we were not able to extensively optimize hyperparameters for this scale. Note that training ViL is still quite expensive to train as there does not exist an optimized hardware implementation yet, which does exist for other models like ViT [15, 14] or Vim [23, 67]. Nevertheless, the efficient alternating block design of ViL makes it already faster than Vim (up to 69% as shown in Appendix B.1) while only using torch.compile, a generic speed optimization of the PyTorch [45] framework.
Model Epochs #Params FLOPS IN-1K
DeiT-T [51] 300 6M 1.3G 72.2
DeiT-II-T [53] 400 6M 1.3G 73.5
DeiT-III-T (reimpl.) 800+20 6M 1.3G 76.2
VRWKV-T [19] 300 6M 1.2G 75.1
Vim-T [67] 300 7M 1.5G 76.1
Mamba®-T [58] 280+20 9M - 77.4
ViL-T 800+20 6M 1.5G 78.3
DeiT-S [51] 300 22M 4.6G 79.8
DeiT-II-S [53] 400 22M 4.6G 80.7
DeiT-III-S [54] 800+20 22M 4.6G 81.4
ConvNeXt-S [41] (iso.) 300 22M 4.3G 79.7
VRWKV-S [19] 300 24M 4.6G 80.1
Vim-S [67] 300 26M 5.3G 80.5
Mamba®-S [58] 280+20 28M - 81.1
ViL-S 400+20 23M 5.1G 81.5
DeiT-B [51] 300 86M 17.6G 81.8
DeiT-II-B [53] 400 86M 17.6G 82.7
DeiT-III-B [54] 800+20 86M 17.6G 83.7
ConvNeXt-B [41] (iso.) 300 87M 16.9G 82.0
VRWKV-B [19] 300 94M 18.2G 82.0
Mamba®-B [58] 280+20 99M - 82.9
ViL-B 400+5 89M 18.6G 82.4
Table 1: ImageNet-1K pre-training accuracy. All models use a patch size of 16x16 with 224x224 resolution at most. Models with “+” in their “Epochs” column pre-train on lower resolution followed by fine-tuning on 224x224 resolution for some epochs. ViL performs favorably against against an isotropic convolutional architecture (ConvNeXt) and vision adaptions of transformers (DeiT), RWKV [47] (VRWKV) and Mamba [23] (Vim, Mamba®).
3.2 ADE20K Semantic Segmentation
Table 2 shows results for transferring ImageNet-1K pre-trained models to ADE20K [66] semantic segmentation using UperNet [61]. Also here, ViL shows strong performances across the board, even outperforming DeiT-III-B despite the lower ImageNet-1K accuracy of ViL-B.
For this benchmark, we train on 512x512 images and inference could be conducted on the original images (without resizing), which results in even larger images. Consequently, the number of patch tokens can become quite large. As sequence length increases, ViL would ideally use a recurrent or chunkwise formulation of the mLSTM layer. Shortly summarized, the mLSTM can operate in parallel, recurrent or chunked modes, each with distinct FLOPS and runtime characteristics. Given a sequence length $L$ and hidden dimension $D$, the trade-off between modes can be formulated as $\mathcal{O}(LCD+LD^{2})$ where $C$ denotes the chunksize of the chunked form. $C=1$ (recurrent mode) yields complexity $\mathcal{O}(LD^{2})$ and $C=L$ (parallel mode) gives $\mathcal{O}(L^{2}D)$ or $\mathcal{O}(L\frac{L}{2}D)$ when omitting the causally masked operations. Due to this trade-off and the current lack of an optimized hardware implementation for the mLSTM, it is difficult to create runtime or FLOPS estimates that are of practical relevance in this setting and therefore leave it to future work.
Single-scale Multi-scale
Model #Params mIoU ACC mIoU ACC
DeiT-T [51] 10M 38.1 78.2 40.3 79.9
DeiT-III-T [54] 10M 39.8 79.2 42.2 80.7
Vim-T [67] 13M 41.0 - - -
ViL-T 11M 41.2 80.2 43.1 81.3
DeiT-S [51] 41M 43.1 80.7 45.2 81.8
DeiT-III-S [54] 41M 45.2 81.5 46.3 82.3
Vim-S [67] 46M 44.9 - - -
Mamba®-S [58] 56M 45.3 - - -
ViL-S 42M 46.3 82.0 47.9 82.9
DeiT-B [51] 113M 45.8 82.1 47.0 82.9
DeiT-III-B [54] 113M 47.5 82.6 49.0 83.3
Mamba®-B [58] 132M 47.7 - - -
ViL-B 115M 48.6 82.8 49.6 83.3
Table 2: Semantic segmentation results on ADE20K [66] using UperNet [61]. We report mean intersection over union (mIoU) and pixelwise accuracy (ACC) for single- and multi-scale evaluation. Models are trained for 160K updates with a batchsize of 16 on 512x512 resolution. Detailed hyperparameters are listed in Appendix Table 14.
3.3 VTAB-1K Transfer Classification
Table 3 shows transfer classification results for ImageNet-1K pre-trained models on the VTAB-1K [64] benchmark. VTAB-1K consists of 19 datasets split into 7 natural datasets (such as CIFAR100 [38], Caltech101 [20] or SUN397 [60]), 4 specialized datasets (medical imaging [57, 34] and remote sensing [28, 11]) and 8 structured datasets (with tasks such as object counting [33] or binned depth estimation [21]). We follow common practices and tune the learning rate per model and dataset on the validation set followed by training with the best learning rate on the union of train and validation set. The performance metric is the average testset accuracy over 5 seeds. ViL shows strong transfer classification performance outperforming all other models on the average over all 19 datasets. ViL performs particularly well on the structured datasets where ViL-B outperforms DeiT-III-B despite ViL-B having lower ImageNet-1K accuracy.
Model #Params FLOPS Natural Specialized Structured Average
DeiT-T [51] 6M 1.3G 69.2 82.0 53.3 65.2
DeiT-III-T (reimpl.) 6M 1.3G 71.9 82.6 55.2 67.1
Vim-T [67] 7M 1.5G 68.0 80.7 47.1 61.9
ViL-T 6M 1.5G 73.6 83.4 56.1 68.3
DeiT-S [51] 22M 4.6G 73.3 83.8 53.2 67.1
DeiT-III-S [54] 22M 4.6G 75.0 83.2 52.3 67.2
Vim-S [67] 26M 5.3G 69.6 81.7 49.4 63.6
ViL-S 23M 5.1G 75.3 84.3 58.3 70.0
DeiT-B [51] 86M 17.6G 76.5 85.2 55.7 69.6
DeiT-III-B [54] 86M 17.6G 77.6 84.8 56.6 70.3
ViL-B 89M 18.6G 76.6 84.7 59.1 70.9
Table 3: Transfer classification accuracies on the VTAB-1K [64] benchmark using ImageNet-1K pre-trained models. VTAB-1K consists of 19 datasets split into 7 natural, 4 specialized and 8 structured datasets. We show averages per category and the average accuracy over all 19 datasets (Appendix Table 10 lists all individual accuracies). ViL shows strong transfer classification performance, outperforming heavily optimized ViT protocols and Vim on the full VTAB-1K benchmark. Notably, ViL performs exceptionally well on the structured category. We tune the learning rate for each model and dataset on the validation set and report the average performance on the test set over 5 seeds. Appendix Table 13 lists further hyperparameters.
3.4 Block Design
@@
Table 4: Traversing the sequence of patch tokens in multiple directions yields better performance than the plain, uni-directional xLSTM block. However, using multiple directions per block also requires more compute and parameters. Sharing parameters between directions matches #Params but overloads them, leading to performance drops. Using multi-directional designs but alternating between them in consecutive blocks, matches #Params and FLOPS while achieving good performances. While quad-directional designs yields the best performances, they comes with drastically increased runtime. As our current quad-directional implementation is not compatible with torch.compile, it suffers from increased runtime even when alternating between directions. Since quad-directional designs are compute intensive, this study was conducted by training a ViL-T on an ImageNet-1K subset of 100 classes using resolution 128. “Param. Sharing” shares the parameters of the two (bidirectional) or four (quaddirectional) mLSTM layers in each block.
@@ 3.5 Classification Design
In order to perform classification with a ViT, the sequence of tokens is pooled into a single token, which is then used as input to a classification head. The most common methods to do this pooling are (i) adding a learnable [CLS] token at the start of the sequence or (ii) averaging all patch tokens to produce an [AVG] token. Whether to use the [CLS] or [AVG] token is typically a hyperparameter, where both variants achieve comparable performances. On the contrary, autoregressive models often require specialized classification designs. For example, Vim [67] requires the [CLS] token to be in the middle of the sequence, suffering heavy performance losses if other classification designs, e.g., an [AVG] token or two [CLS] tokens at start and end of the sequence, are employed. Due to its autoregressive nature, we explore different classification designs for ViL in Table 5. [AVG] averages all patch tokens, “Middle Patch” uses the middle patch token, “Middle [CLS]” uses a [CLS] token in the middle of the sequence, “Bilateral [AVG]” uses the average of the first and last patch token. We find ViL to be relatively robust the classification design where all performances that use a single token are within 0.6%. We only add a [CLS] token if it is used for classification, hence the best design (“Bilateral Concat”) does not use a [CLS] token.
Design Dim ACC
[AVG] 192 75.4
Middle Patch 192 75.9
Middle [CLS] 192 75.3
Bilateral [AVG] 192 75.9
Bilateral Concat 384 76.4
@@
Table 5: How to pool the sequence of tokens for classification? ViL is relatively robust to the classification design. Using the middle patch token and averaging the first and last token (“Bilateral [AVG]”) work similarly well. The good performance of middle patch could be caused by the bias in ImageNet-1K that objects are typically in the center of the image. We therefore choose a bilateral design to avoid exploiting this bias. Additionally, concatenating the first and last token (“Bilateral Concat”) instead of averaging improves performance and is used as final design for ViL. We conduct this study by training a ViL-T on ImageNet-1K for 400 epochs.
3.6 Vision Adjustments of xLSTM
The xLSTM [5] architecture was designed for language modeling and therefore includes some design choices that are specific to language modeling. One of these design choices is a causal 1D convolution (with kernel size 4) to encode local context into the query and key input of the mLSTM layer. As images are neither causal nor 1D structures, we replace the causal 1D convolution with a 2D convolution (with kernel size 3). When training a ViL-T for 400 epochs, this improves ImageNet-1K accuracy from 76.1% to 76.7%.
Another language specific design choice is the absence of bias parameters in projection and layernorm [4] layers which improves training stability for larger models [12, 16]. As vision models are typically much smaller than language models, ViTs commonly use all bias parameters. When adding bias parameters, a ViL-T trained for 400 epochs improves from 76.7% to 77.1% on ImageNet-1K.
@@ 4 Conclusion
Motivated by the success of xLSTM in language modeling, we introduced ViL, an adaption of the xLSTM architecture to vision tasks. ViL processes a sequence of patch tokens in alternating fashion. Odd blocks process image patches row-wise from top left to bottom right and even blocks go from bottom right to top left. Our new architecture outperforms SSM-based vision architectures and also optimized ViT models on ImageNet-1K classification, VTAB-1K transfer classification and ADE20K semantic segmentation. Remarkably, ViL is able to outperform ViT training pipelines, which are the result of years of hyperparameter tuning and transformer improvements.
In the future, we see potential in applying ViL when high-resolution images are needed for optimal performance, such as semantic segmentation or medical imaging. In these settings, transformers suffer from high computational costs due to the quadratic complexity of self-attention, where ViL can use a chunked form to trade-off between its parallel form (quadratic complexity) and recurrent form (linear complexity). Additionally, improving pre-training schemes (e.g. via self-supervised learning), exploring better hyperparameter settings or investigating hierarchical architectures are promising future directions that could improve ViL even further.
@@ Acknowledgments
Appendix A Changelog
A.1 Version 1
In the first version of this paper, we trained models with a causal 1D convolution in the mLSTM layer (instead of a 2D convolution), “Bilateral [AVG]” pooling (instead of “Bilateral Concat”) and a constant resolution of 224 (instead of pre-training at 192 followed by a short fine-tuning on resolution 224). This version of the architecture is also used to conduct the studies in Table 4 and Table 5. For completeness, we report the accuracies of these models in Table 6. In the first version, we trained slightly deeper models (ViL-T+ with 29 blocks and ViL-S+ with 26 blocks) to match the parameter count of Vim-T and Vim-S. However, as the models with a regular depth of 24 were already better than Vim with less parameters, we omit them in future versions.
Method #Params ACC
ViL-T 6M 77.3
ViL-T+ 7M 78.1
ViL-S 23M 80.7
ViL-S+ 26M 80.9
ViL-B 88M 81.6
Model #Params ACC
Long-sequence fine-tuning
ViL-T+ 7M 80.0
ViL-S+ 26M 82.2
ViL-B 88M 82.7
Table 6: ImageNet-1K accuracies of models trained in the first version of this paper.
A.2 Version 2
The current version of the paper improves performance by (i) replacing the causal 1D convolution with a 2D convolution (with kernel size 3 and padding 1) (ii) using “Bilateral Concat” instead of “Bilateral [AVG]” pooling (iii) using bias parameters in projection and LayerNorm [4] layers (iv) pre-training models on 192x192 followed by a short fine-tuning on 224x224 resolution.
Appendix B Extended Results
B.1 Runtime Comparison of ViL vs Vim